in a two digit number,if one is added to the ten's digit,it is four times the unit's digit.if 18 times the unit's digit is subtracted from the number then the number becomes half it's original value.find the two digit number
Answer by ankor@dixie-net.com(22740)

in a two digit number,if one is added to the ten's digit,it is four times the unit's digit.
if 18 times the unit's digit is subtracted from the number then the number
becomes half it's original value.
find the two digit number
:
Let x = the 10's digit
Let y = the units
then
10x+y = the number
:
Write an equation for each statement:
:
"if one is added to the ten's digit, it is four times the unit's digit."
x + 1 = 4y
x = 4y - 1
:
"if 18 times the unit's digit is subtracted from the number then the number
becomes half it's original value."
10x+y - 18y = .5(10x+y)
10x - 17y = 5x + .5y
10x - 5x = .5y + 17y
5x = 17.5y
Replace x with (4y-1)
5(4y-1) = 17.5y
20y - 5 = 17.5y
20y - 17.5y = 5
2.5y = 5
y = 5/2.5
y = 2
find x
x = 4y - 1
x = 7
:
72 is the number
:
:
Check solution in the statement:
if 18 times the unit's digit is subtracted from the number then the number becomes half it's original value."
72 - 18(2) = .5(72)
72 - 36 = 36, confirms our solution